OpenStudy (anonymous):

Chris can be paid in one of two ways. Plan "A" is a salary of $350 per month, plus a commission of 7% of sales. Plan "B" is a salary of $630 per month, plus a commission of 3% of sales. Chris should select plan "A" for sales greater than?

OpenStudy (anonymous):

$7000 Correct?

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

do you have option

OpenStudy (gorv):

let sale = x

OpenStudy (anonymous):

No options :/ It's like a free response one.

OpenStudy (gorv):

now calculate the payment he will get

OpenStudy (anonymous):

Yes, I put in a random number for x and narrowed it down to 7000 dollars

OpenStudy (gorv):

\[350+x *\frac{ 7 }{ 100 }\]

OpenStudy (gorv):

that one for A

OpenStudy (anonymous):

So B i: 630 + x * .03

OpenStudy (anonymous):

B is*

OpenStudy (gorv):

\[630+x*\frac{ 3 }{ 100 }\]

OpenStudy (anonymous):

Yes, so if x is $7000, then Chris should pick plan "A" when sales are greater that: $7000 correct?

OpenStudy (gorv):

so for this \[350+\frac{ 7x }{ 100}>630+\frac{ 3x }{ 100 }\]

OpenStudy (gorv):

subtract 350 from both side

OpenStudy (anonymous):

\[7x \div100 > 280 + 3x \div100\]

OpenStudy (gorv):

now multiply by 100 both side

OpenStudy (anonymous):

\[700x \div100 > 3x\]

OpenStudy (anonymous):

7x>3x

OpenStudy (gorv):

loll 280 is also there

OpenStudy (anonymous):

Crap, lol hold up.

OpenStudy (gorv):

\[\frac{ 7x }{ 100 }*100>280*100+\frac{ 3x }{ 100 }*100\]

OpenStudy (gorv):

now give me equation after cancelling 100

OpenStudy (anonymous):

7x> 28000 + 3x

OpenStudy (anonymous):

4x > 28000

OpenStudy (anonymous):

x > 7000

OpenStudy (gorv):

yepppppppppppp

OpenStudy (anonymous):

Awesome. Thank you very much @gorv